OpenStudy (anonymous):

Evaluate the given expressions. All numbers are approximate. 1328^2-(-7.023)^4/1.037(-2.01) how do i do this?

OpenStudy (anonymous):

they have = -2256 at top and -2.08437 at bottom and this gives you 1082.341427 wich equalls 0ff to 1082. how did they come up with -2256

OpenStudy (anonymous):

\[ 1328^2-(-7.023)^4 \over 1.037(-2.01) \] \[ =\frac{1763584-2432.711866955841}{(- 2.08437)}\] \[ =\frac{1761151.288133044159}{(- 2.08437)}=-844932.18005106778499\] \[=-844932.18005106778499 \approx -844932.18\]

OpenStudy (anonymous):

some how they have -2256/-2.08437=1082.341427=1080

OpenStudy (anonymous):

how did they get -2256